# On the metric of the jet bundle and similarity on weighted Dirichlet spaces

### v2

@@ 4. Similarity on weighted Dirichlet spaces
Is the ratio of the metrics of holomorphic line bundles a similarity invariant of the backward shift operator on the weighted Dirichlet space?
The similarity invariant of backward shift operators on weighted Dirichlet spaces cannot be obtained from the previous statements about Hardy space and weighted Bergman spaces. The main reason weighted Dirichlet spaces differ from Hardy space and weighted Bergman spaces in Question 1 is that the multiplier algebra of weighted Dirichlet spaces is not equal to $H^{\infty}(\mathbb{D})$. For the similarity of backward shift operators on weighted Dirichlet spaces, assuming that the ratio of the metrics is bounded and bounded from 00, we introduce new tools such as holomorphic jet bundles and provide two sufficient conditions in Theorem 4.3 and Theorem 4.9, respectively. Meanwhile, as shown in Example 4.8 and Example 4.10, it is challenging to construct a backward shift operator that satisfies the following two conditions: (1) the ratio of the metric of the holomorphic line bundle induced by the operator to the metric of the holomorphic line bundle associated with the backward shift operator on Drichilet space is bounded (or unbounded or tends to 00 as $|w|\rightarrow 1$); (2) the operator is not similar to the backward shift operator on Drichilet space.
@@ Acknowledgment
We gratefully acknowledge Y. H. Zhang and J. M. Yang for their valuable input and many helpful conversations regarding this paper.

### v3

@@ 4. Similarity on weighted Dirichlet spaces
Is the ratio of the metrics of holomorphic line bundles a similarity invariant for the backward shift operator on the weighted Dirichlet space?
A similarity invariant for backward shift operators on weighted Dirichlet spaces, unfortunately, cannot be obtained from previous statements about Hardy and weighted Bergman spaces. The main reason for this has to do with the corresponding multiplier algebra; as is well-known, the multiplier algebra of weighted Dirichlet spaces is not equal to $H^{\infty}(\mathbb{D})$. For the similarity of backward shift operators on weighted Dirichlet spaces, assuming that the ratio of the metrics is bounded and bounded away from 00, we have introduced new perspective on the problem by considering holomorphic jet bundles and provided two sufficient conditions in Theorems 4.3 and 4.9. Meanwhile, as can be seen through Examples 4.8 and 4.11, it is challenging to construct a backward shift operator that satisfies the following two conditions: (1) the ratio of the metric of the holomorphic line bundle induced by the operator to the metric of the holomorphic line bundle associated with the backward shift operator on the Dirichlet space is bounded (or unbounded or tends to 00 as $|w|\rightarrow 1$), (2) the operator is not similar to the backward shift operator on the Dirichlet space.
@@ Acknowledgment
We gratefully acknowledge the valuable help of Y. H. Zhang and J. M. Yang on Example 4.8.